THE SUM OF MR LEE'S AND HIS SON'S AGE IS 49. IN THREE YEARS TIME, HE WILL BE 4 TIMES AS OLD AS HIS SON. FIND THEIR PRESENT AGES
Answer by Alan3354(69443)

THE SUM OF MR LEE'S AND HIS SON'S AGE IS 49. IN THREE YEARS TIME, HE WILL BE 4 TIMES AS OLD AS HIS SON. FIND THEIR PRESENT AGES
-----------
L = Mr. Lee's age
S = Son's age
----------
S+L = 49
L+3 = 4*(S+3)
----------
Sub for L, L = 49-S
49-S + 3 = 4*(S+3)
52-S = 4S+12
40 = 5S
S = 8
----
L = 49-S
L = 41